Arnell Industries has $10 million in permanent debt outstanding. The firm will pay interest only on this debt. Arnell’s marginal tax rate is expected to be 35% for the foreseeable future.

A. Suppose Arnell pays interest of 6% per year on its debt. What is its annual interest tax shield?

B. What is the present value of the interest tax shield, assuming its risk is the same as the loan?

C. Suppose instead that the interest rate on the debt is 5%. What is the present value of the interest tax shield in this case?
Answer #1

A

Tax shield = debt*tax rate*interest rate = 10*0.35*0.06 = 0.21 m

B

PV = tax shield/interest rate = 0.21/0.06 = 3.5m

C

PV = debt*tax rate = 10*0.35 = 3.5m